# Chapter 4 Discrete subgroups of semisimple Lie groups, beyond lattices

### 26 Feb 2024 v1

@@ 3.2.1 Ping pong in projective space
In particular, if we fix $\varepsilon>0$, then any large power of $g$ sends the complement of the open uniform $\varepsilon$-neighbourhood $B_{g^{-1}}^{\varepsilon}$ of $X_{g}^{-}$ into the closure of the open ball $b_{g}^{\varepsilon}$ of radius $\varepsilon$ centred at $x_{g}^{+}$ for $\mathtt{d}_{\mathbb{P}(\mathbb{R}^{d})}$, and $g^{-1}$ has a similar behavior after replacing $(B_{g^{-1}}^{\varepsilon},X_{g}^{-},b_{g}^{\varepsilon},x_{g}^{+})$ by $(B_{g}^{\varepsilon},X_{g}^{+},b_{g^{-1}}^{\varepsilon},x_{g}^{-})$ (see Figure 4).
@@
Similarly to the classical strong Schottky groups of Section 3.1, the group $\Gamma$ admits nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma,G)$, obtained by independently deforming the image of each generator $\gamma_{i}$; moreover, there is a neighbourhood of the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ consisting entirely of injective and discrete representations.
3.2.2 Schottky groups with disjoint ping pong domains
In certain situations it is possible to construct discrete subgroups of $G$, with ping pong dynamics, for which the ping pong domains are pairwise disjoint, as in the case of the classical rank-one Schottky groups of Section 3.1. Achieving this disjointness may require using a slightly modified ping pong compared to Figures 4 and 5, allowing the attracting and repelling subsets of the generators to be larger than points.
Such a construction has been made in $G=\mathrm{PGL}(2n,\mathbb{K})$, acting on the projective space $\mathbb{P}(\mathbb{K}^{2n})$, for $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$: the first examples were constructed by Nori in the 1980s, for $\mathbb{K}=\mathbb{C}$, then generalised by Seade and Verjovsky (see [122]); it was observed in [89] that the construction also works for $\mathbb{K}=\mathbb{R}$. The idea is to consider pairwise disjoint $(n-1)$-dimensional projective subspaces $X_{1}^{+},X_{1}^{-},\dots,X_{m}^{+},X_{m}^{-}$ of $\mathbb{P}(\mathbb{K}^{2n})$ and elements $\gamma_{1},\dots,\gamma_{m}\in G$ such that for any $1\leq i\leq m$ we have $\gamma_{i}^{\pm k}\cdot x\to X_{i}^{\pm}$ for all $x\in\mathbb{P}(\mathbb{K}^{2n})\smallsetminus X_{i}^{\mp}$ as $k\to+\infty$, uniformly on compact sets. Up to replacing each $\gamma_{i}$ by a large power, we may assume that there exist tubular neighbourhoods $B_{i}^{\pm}$ of $X_{i}^{\pm}$ such that $B_{1}^{+},B_{1}^{-},\dots,B_{m}^{+},B_{m}^{-}$ are pairwise disjoint, $\mathbb{P}(\mathbb{K}^{2n})\smallsetminus\bigcup_{i=1}^{m}(B_{i}^{-}\cup B_{i}^{+})$ has nonempty interior, and $\gamma_{i}\cdot\mathrm{Int}(\mathbb{P}(\mathbb{K}^{2n})\smallsetminus B_{i}^{-})=B_{i}^{+}$ for all $i$. Then the subgroup $\Gamma$ of $G$ generated by $\gamma_{1},\dots,\gamma_{m}$ is a nonabelian free group with free generating subset $\{\gamma_{1},\dots,\gamma_{m}\}$. It is discrete in $G$, and it acts properly discontinuously on $\Omega:=\mathrm{Int}(\bigcup_{\gamma\in\Gamma}\gamma\cdot\nolinebreak\mathcal{D})$ with fundamental domain $\mathcal{D}$, as in Remark 4.7. As for the classical strong Schottky groups of Section 3.1, there is a neighbourhood of the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ consisting entirely of injective and discrete representations.
3.2.3 Crooked Schottky groups
@@
Let $G=\mathrm{Sp}(2n,\mathbb{R})$ be the group of elements of $\mathrm{GL}(2n,\mathbb{R})$ that preserve the skew-symmetric bilinear form $\omega(v,v^{\prime})=\sum_{i=1}^{2n}(-1)^{i}v_{i}v^{\prime}_{2n+1-i}$ on $\mathbb{R}^{2n}$. A symplectic basis of $\mathbb{R}^{2n}$ is a basis in which the matrix of $\omega$ is antidiagonal with entries $1,-1,\dots,1,-1$; for instance, the canonical basis is a symplectic basis. To any symplectic basis $(e_{1},\dots,e_{2n})$ of $\mathbb{R}^{2n}$ we associate an open simplex $B=\mathbb{P}(\mathbb{R}^{>0}\text{-span}(e_{1},\dots,e_{2n}))$ in $\mathbb{P}(\mathbb{R}^{2n})$, which we call a symplectic simplex. Its dual $B^{*}:=\{[v]\in\nolinebreak\mathbb{P}(\mathbb{R}^{2n})\,|\,\mathbb{P}(v^{\perp})\cap\nolinebreak\overline{B}=\nolinebreak\emptyset\}$ (where $v^{\perp}$ denotes the orthogonal of $v$ with respect to $\omega$ and $\overline{B}$ the closure of $B$ in $\mathbb{P}(\mathbb{R}^{2n})$) is still a symplectic simplex, associated to the symplectic basis $(e_{2n},-e_{2n-1},\dots,e_{2},-e_{1})$. Note that $B$ and $B^{*}$ are two of the $2^{2n-1}$ connected components of $\mathbb{P}(\mathbb{R}^{2n})\smallsetminus\bigcup_{i=1}^{2n}e_{i}^{\perp}$ (see Figure 6). We also make the elementary observation that for any symplectic simplices $B_{1}$ and $B_{2}$, we have $B_{1}\subset B_{2}^{*}$ if and only if $B_{2}\subset B_{1}^{*}$.
@@
As in the classical case of Section 3.1, continuous deformations of configurations of symplectic simplices yield nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma,G)$ of crooked Schottky groups $\Gamma$ for which each $\rho_{t}$ is injective with discrete image. If the symplectic simplices $B_{1}^{\pm},\dots,B_{m}^{\pm}$ in the initial configuration have pairwise disjoint closures (i.e. $\Gamma$ is a “strong” crooked Schottky group), then the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ admits an open neighbourhood in $\mathrm{Hom}(\Gamma,G)$ consisting entirely of injective and discrete representations.
3.3 Higher-rank deformations of Fuchsian representations
Inspired by the quasi-Fuchsian representations and their higher-dimensional analogues from Section 3.1, here is one strategy for constructing “flexible” infinite discrete subgroups, beyond nonabelian free groups, in semisimple Lie groups $G$ of higher real rank. Consider a finitely generated group $\Gamma_{0}$, an injective and discrete representation $\sigma_{0}$ of $\Gamma_{0}$ into a simple Lie group $G^{\prime}$ of real rank one, and a nontrivial Lie group homomorphism $\tau:G^{\prime}\to G$. Consider the composed representation
@@ 3.3 Higher-rank deformations of Fuchsian representations
In some important cases (see e.g. Facts 4.30 and 4.31), there will be an open neighbourhood of $\rho_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations. The goal is then to deform $\rho_{0}$ nontrivially in $\mathrm{Hom}(\Gamma_{0},G)$ outside of $\mathrm{Hom}(\Gamma_{0},G^{\prime})$, so as to obtain discrete subgroups of $G$ that are isomorphic to $\Gamma_{0}$ but not conjugate to $\Gamma_{0}$ or any subgroup of $G^{\prime}$ (ideally Zariski-dense discrete subgroups of $G$).
This strategy works well, for instance, for $\Gamma_{0}=\pi_{1}(S)$ where $S$ is a closed orientable surface of genus $\geq 2$ as in Section 3.1, and $G^{\prime}=\mathrm{SL}(2,\mathbb{R})$ or $\mathrm{PSL}(2,\mathbb{R})$. Let us give three examples in this setting.
3.3.1 Barbot representations
For $d\geq 2$, consider the standard embedding $\tau:G^{\prime}=\mathrm{SL}(2,\mathbb{R})\hookrightarrow G=\mathrm{SL}(d,\mathbb{R})$, preserving a $2$-dimensional linear subspace of $\mathbb{R}^{d}$. Then there is a neighbourhood of $\rho_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations (see Section 5.1). Nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma_{0},G)$ exist; for $d=3$, they were studied by Barbot, who particularly investigated [8] the case that the $\rho_{t}$ take values in $\mathrm{GL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$, seen as the subgroup of $G$ consisting of lower block-triangular matrices with blocks of size $(2,1)$.
3.3.2 Hitchin representations
@@
The proofs of Labourie and Fock–Goncharov for general $d$ are dynamical. They involve two key objects. The first one is the Gromov boundary $\partial_{\infty}\Gamma_{0}$ of $\Gamma_{0}$: by definition, this is the visual boundary of a proper geodesic metric space on which $\Gamma_{0}$ acts properly discontinuously, by isometries, with compact quotient; in our situation, $\Gamma_{0}$ is a closed surface group and $\partial_{\infty}\Gamma_{0}$ is the visual boundary of $\mathbb{H}^{2}$, namely a circle. The second key object is the space $\mathrm{Flags}(\mathbb{R}^{d})$ of full flags $(V_{1}\subset\dots\subset V_{d-1}\subset\mathbb{R}^{d})$ of $\mathbb{R}^{d}$ (where each $V_{i}$ is an $i$-dimensional linear subspace of $\mathbb{R}^{d}$); this space $\mathrm{Flags}(\mathbb{R}^{d})$ is compact with a transitive action of $G=\mathrm{PSL}(d,\mathbb{R})$, and may be thought of as a kind of “boundary” for $G$ or its symmetric space. The point of the proof is then to show that for any Hitchin representation $\rho:\Gamma_{0}\to G$, there exists a continuous, injective, $\rho$-equivariant “boundary map” $\xi:\partial_{\infty}\Gamma_{0}\to\mathrm{Flags}(\mathbb{R}^{d})$. See Figure 10. (By $\rho$-equivariant we mean that $\xi(\gamma\cdot z)=\rho(\gamma)\cdot\xi(z)$ for all $\gamma\in\Gamma_{0}$ and all $z\in\partial_{\infty}\Gamma_{0}$.) The existence of such a boundary map $\xi$ easily implies that $\rho$ is injective and discrete: see the proof of Lemma 4.27.(3). Indeed, the idea is that the continuous, injective, equivariant map $\xi$ “transfers”, to $\mathrm{Flags}(\mathbb{R}^{d})$, the dynamics of the intrinsic action of $\Gamma_{0}$ on $\partial_{\infty}\Gamma_{0}$, which is a so-called convergence action: any sequence $(\gamma_{k})_{k\in\mathbb{N}}$ of pairwise distinct elements of $\Gamma_{0}$ comes with some contraction in $\partial_{\infty}\Gamma_{0}$, hence (using $\xi$) the sequence $(\rho(\gamma_{k}))_{k\in\mathbb{N}}$ comes with some contraction in $\mathrm{Flags}(\mathbb{R}^{d})$, and this prevents $(\rho(\gamma_{k}))_{k\in\mathbb{N}}$ from converging to the identity element of $G$.
@@
3.3.3 Maximal representations
For $n\geq 2$, consider the embedding $\tau:G^{\prime}=\mathrm{PSL}(2,\mathbb{R})\simeq\mathrm{SO}(2,1)_{0}\hookrightarrow\mathrm{SO}(2,n)$. Then the entire connected component of $\rho_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consists of injective and discrete representations, as was proved by Burger, Iozzi, and Wienhard [37]. This is an example of a so-called maximal component: it consists of representations (called maximal representations) that maximise the Toledo invariant (a topological invariant generalising the Euler number, see e.g. [38, § 5.1]).
3.4 Higher Teichmüller theory
We already encountered in Sections 2.4 and 3.1 the Teichmüller space of a closed surface $S$ of genus $\geq 2$. It is a fundamental object in many areas of mathematics, which can be viewed both as a moduli space for marked complex structures on $S$ or, via the Uniformisation Theorem, as a moduli space for marked hyperbolic structures on $S$. In this second point of view, the holonomy representation of the fundamental group $\Gamma_{0}=\pi_{1}(S)$ naturally realises the Teichmüller space of $S$ as a connected component of the $G$-character variety of $\Gamma_{0}$ for $G=\mathrm{PSL}(2,\mathbb{R})$, corresponding to the image, modulo conjugation by $G$ at the target, of a connected component of $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations.
An interesting and perhaps surprising phenomenon, which has led to a considerable amount of research in the past twenty years, is that for certain semisimple Lie groups $G$ of higher real rank, there also exist connected components of $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations, and which are nontrivial in the sense that they are not reduced to a single representation and its conjugates by $G$. The images in the $G$-character variety of these components are now called higher(-rank) Teichmüller spaces. We saw two examples in Section 3.3:

### 28 Mar 2024 v2

@@ 2 Ping pong in higher real rank
In particular, if we fix $\varepsilon>0$, then any large power of $g$ sends the complement of the open uniform $\varepsilon$-neighbourhood $B_{g^{-1}}^{\varepsilon}$ of $X_{g}^{-}$ into the closure of the open ball $b_{g}^{\varepsilon}$ of radius $\varepsilon$ centred at $x_{g}^{+}$ for $\mathtt{d}_{\mathbb{P}(\mathbb{R}^{d})}$, and $g^{-1}$ has a similar behaviour after replacing $(B_{g^{-1}}^{\varepsilon},X_{g}^{-},b_{g}^{\varepsilon},x_{g}^{+})$ by $(B_{g}^{\varepsilon},X_{g}^{+},b_{g^{-1}}^{\varepsilon},x_{g}^{-})$ (see Figure 4).
@@
Similarly to the classical strong Schottky groups of Section 1, the group $\Gamma$ admits nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma,G)$, obtained by independently deforming the image of each generator $\gamma_{i}$; moreover, there is a neighbourhood of the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ consisting entirely of injective and discrete representations.
Schottky groups with disjoint ping pong domains
In certain situations it is possible to construct discrete subgroups of $G$, with ping pong dynamics, for which the ping pong domains are pairwise disjoint, as in the case of the classical rank-one Schottky groups of Section 1. Achieving this disjointness may require using a slightly modified ping pong compared to Figures 4 and 5, allowing the attracting and repelling subsets of the generators to be larger than points.
Such a construction has been made in $G=\mathrm{PGL}(2n,\mathbb{K})$, acting on the projective space $\mathbb{P}(\mathbb{K}^{2n})$, for $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$: the first examples were constructed by Nori in the 1980s, for $\mathbb{K}=\mathbb{C}$, then generalised by Seade and Verjovsky (see [124]); it was observed in [91] that the construction also works for $\mathbb{K}=\mathbb{R}$. The idea is to consider pairwise disjoint $(n-1)$-dimensional projective subspaces $X_{1}^{+},X_{1}^{-},\dots,X_{m}^{+},X_{m}^{-}$ of $\mathbb{P}(\mathbb{K}^{2n})$ and elements $\gamma_{1},\dots,\gamma_{m}\in G$ such that for any $1\leq i\leq m$ we have $\gamma_{i}^{\pm k}\cdot x\to X_{i}^{\pm}$ for all $x\in\mathbb{P}(\mathbb{K}^{2n})\smallsetminus X_{i}^{\mp}$ as $k\to+\infty$, uniformly on compact sets. Up to replacing each $\gamma_{i}$ by a large power, we may assume that there exist tubular neighbourhoods $B_{i}^{\pm}$ of $X_{i}^{\pm}$ such that $B_{1}^{+},B_{1}^{-},\dots,B_{m}^{+},B_{m}^{-}$ are pairwise disjoint, $\mathbb{P}(\mathbb{K}^{2n})\smallsetminus\bigcup_{i=1}^{m}(B_{i}^{-}\cup B_{i}^{+})$ has nonempty interior, and $\gamma_{i}\cdot\mathrm{Int}(\mathbb{P}(\mathbb{K}^{2n})\smallsetminus B_{i}^{-})=B_{i}^{+}$ for all $i$. Then the subgroup $\Gamma$ of $G$ generated by $\gamma_{1},\dots,\gamma_{m}$ is a nonabelian free group with free generating subset $\{\gamma_{1},\dots,\gamma_{m}\}$. It is discrete in $G$, and it acts properly discontinuously on $\Omega:=\mathrm{Int}(\bigcup_{\gamma\in\Gamma}\gamma\cdot\nolinebreak\mathcal{D})$ with fundamental domain $\mathcal{D}$, as in Remark 3.3. As for the classical strong Schottky groups of Section 1, there is a neighbourhood of the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ consisting entirely of injective and discrete representations.
Crooked Schottky groups
@@
Let $G=\mathrm{Sp}(2n,\mathbb{R})$ be the group of elements of $\mathrm{GL}(2n,\mathbb{R})$ that preserve the skew-symmetric bilinear form $\omega(v,v^{\prime})=\sum_{i=1}^{2n}(-1)^{i}v_{i}v^{\prime}_{2n+1-i}$ on $\mathbb{R}^{2n}$. A symplectic basis of $\mathbb{R}^{2n}$ is a basis in which the matrix of $\omega$ is antidiagonal with entries $1,-1,\dots,1,-1$; for instance, the canonical basis is a symplectic basis. To any symplectic basis $(e_{1},\dots,e_{2n})$ of $\mathbb{R}^{2n}$ we associate an open simplex $B=\mathbb{P}(\mathbb{R}^{>0}\text{-span}(e_{1},\dots,e_{2n}))$ in $\mathbb{P}(\mathbb{R}^{2n})$, which we call a symplectic simplex. Its dual $B^{*}:=\{[v]\in\nolinebreak\mathbb{P}(\mathbb{R}^{2n})\,|\,\mathbb{P}(v^{\perp})\cap\nolinebreak\overline{B}=\nolinebreak\emptyset\}$ (where $v^{\perp}$ denotes the orthogonal of $v$ with respect to $\omega$ and $\overline{B}$ the closure of $B$ in $\mathbb{P}(\mathbb{R}^{2n})$) is still a symplectic simplex, associated to the symplectic basis $(e_{2n},-e_{2n-1},\dots,e_{2},-e_{1})$. Note that $B$ and $B^{*}$ are two of the $2^{2n-1}$ connected components of $\mathbb{P}(\mathbb{R}^{2n})\smallsetminus\bigcup_{i=1}^{2n}\mathbb{P}(e_{i}^{\perp})$ (see Figure 6). We also make the elementary observation that for any symplectic simplices $B_{1}$ and $B_{2}$, we have $B_{1}\subset B_{2}^{*}$ if and only if $B_{2}\subset B_{1}^{*}$.
@@
As in the classical case of Section 1, continuous deformations of configurations of symplectic simplices yield nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma,G)$ of crooked Schottky groups $\Gamma$ for which each $\rho_{t}$ is injective with discrete image. If the symplectic simplices $B_{1}^{\pm},\dots,B_{m}^{\pm}$ in the initial configuration have pairwise disjoint closures (i.e. $\Gamma$ is a “strong” crooked Schottky group), then the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ admits an open neighbourhood in $\mathrm{Hom}(\Gamma,G)$ consisting entirely of injective and discrete representations.
3 Higher-rank deformations of Fuchsian representations
Inspired by the quasi-Fuchsian representations and their higher-dimensional analogues from Section 1, here is one strategy for constructing “flexible” infinite discrete subgroups, beyond nonabelian free groups, in semisimple Lie groups $G$ of higher real rank. Consider a finitely generated group $\Gamma_{0}$, an injective and discrete representation $\sigma_{0}$ of $\Gamma_{0}$ into a simple Lie group $G^{\prime}$ of real rank one, and a nontrivial Lie group homomorphism $\tau:G^{\prime}\to G$. Consider the composed representation
@@ 3 Higher-rank deformations of Fuchsian representations
In some important cases (see e.g. Facts 5.6 and 5.7), there will be an open neighbourhood of $\rho_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations. The goal is then to deform $\rho_{0}$ nontrivially in $\mathrm{Hom}(\Gamma_{0},G)$ outside of $\mathrm{Hom}(\Gamma_{0},G^{\prime})$, so as to obtain discrete subgroups of $G$ that are isomorphic to $\Gamma_{0}$ but not conjugate to $\Gamma_{0}$ or any subgroup of $G^{\prime}$ (ideally Zariski-dense discrete subgroups of $G$).
This strategy works well, for instance, for $\Gamma_{0}=\pi_{1}(S)$ where $S$ is a closed orientable surface of genus $\geq 2$ as in Section 1, and $G^{\prime}=\mathrm{SL}(2,\mathbb{R})$ or $\mathrm{PSL}(2,\mathbb{R})$. Let us give three examples in this setting.
Barbot representations
For $d\geq 2$, consider the standard embedding $\tau:G^{\prime}=\mathrm{SL}(2,\mathbb{R})\hookrightarrow G=\mathrm{SL}(d,\mathbb{R})$, acting trivially on a $(d-2)$-dimensional linear subspace of $\mathbb{R}^{d}$. Then there is a neighbourhood of $\rho_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations (see Section 1). Nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma_{0},G)$ exist; for $d=3$, they were studied by Barbot, who particularly investigated [8] the case that the $\rho_{t}$ take values in $\mathrm{GL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$, seen as the subgroup of $G$ consisting of lower block-triangular matrices with blocks of size $(2,1)$.
Hitchin representations
@@
The proofs of Labourie and Fock–Goncharov for general $d$ are dynamical. They involve two key objects. The first one is the Gromov boundary $\partial_{\infty}\Gamma_{0}$ of $\Gamma_{0}$: by definition, this is the visual boundary of a proper geodesic metric space on which $\Gamma_{0}$ acts properly discontinuously, by isometries, with compact quotient; in our situation, $\Gamma_{0}$ is a closed surface group and $\partial_{\infty}\Gamma_{0}$ is the visual boundary of $\mathbb{H}^{2}$, namely a circle. The second key object is the space $\mathrm{Flags}(\mathbb{R}^{d})$ of full flags $(V_{1}\subset\dots\subset V_{d-1}\subset\mathbb{R}^{d})$ of $\mathbb{R}^{d}$ (where each $V_{i}$ is an $i$-dimensional linear subspace of $\mathbb{R}^{d}$); this space $\mathrm{Flags}(\mathbb{R}^{d})$ is compact with a transitive action of $G=\mathrm{PSL}(d,\mathbb{R})$, and may be thought of as a kind of “boundary” for $G$ or its symmetric space. The point of the proof is then to show that for any Hitchin representation $\rho:\Gamma_{0}\to G$, there exists a continuous, injective, $\rho$-equivariant “boundary map” $\xi:\partial_{\infty}\Gamma_{0}\to\mathrm{Flags}(\mathbb{R}^{d})$. See Figure 10. (By $\rho$-equivariant we mean that $\xi(\gamma\cdot w)=\rho(\gamma)\cdot\xi(w)$ for all $\gamma\in\Gamma_{0}$ and all $w\in\partial_{\infty}\Gamma_{0}$.) The existence of such a boundary map $\xi$ easily implies that $\rho$ is injective and discrete: see the proof of Lemma 5.3.(3). Indeed, the idea is that the continuous, injective, equivariant map $\xi$ “transfers”, to $\mathrm{Flags}(\mathbb{R}^{d})$, the dynamics of the intrinsic action of $\Gamma_{0}$ on $\partial_{\infty}\Gamma_{0}$, which is a so-called convergence action: any sequence $(\gamma_{k})_{k\in\mathbb{N}}$ of pairwise distinct elements of $\Gamma_{0}$ comes with some contraction in $\partial_{\infty}\Gamma_{0}$, hence (using $\xi$) the sequence $(\rho(\gamma_{k}))_{k\in\mathbb{N}}$ comes with some contraction in $\mathrm{Flags}(\mathbb{R}^{d})$, and this prevents $(\rho(\gamma_{k}))_{k\in\mathbb{N}}$ from converging to the identity element of $G$.
@@
Maximal representations
For $n\geq 2$, consider the embedding $\tau:G^{\prime}=\mathrm{PSL}(2,\mathbb{R})\simeq\mathrm{SO}(2,1)_{0}\hookrightarrow\mathrm{SO}(2,n)$. Then the entire connected component of $\rho_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consists of injective and discrete representations, as was proved by Burger, Iozzi, and Wienhard [38]. This is an example of a so-called maximal component: it consists of representations (called maximal representations) that maximise the Toledo invariant (a topological invariant generalising the Euler number, see e.g. [39, § 5.1]).
4 Higher Teichmüller theory
We already encountered in Sections 4 and 1 the Teichmüller space of a closed surface $S$ of genus $\geq 2$. It is a fundamental object in many areas of mathematics, which can be viewed both as a moduli space for marked complex structures on $S$ or, via the Uniformisation Theorem, as a moduli space for marked hyperbolic structures on $S$. In this second point of view, the holonomy representation of the fundamental group $\Gamma_{0}=\pi_{1}(S)$ naturally realises the Teichmüller space of $S$ as a connected component of the $G$-character variety of $\Gamma_{0}$ for $G=\mathrm{PSL}(2,\mathbb{R})$, corresponding to the image, modulo conjugation by $G$ at the target, of a connected component of $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations.
An interesting and perhaps surprising phenomenon, which has led to a considerable amount of research in the past twenty years, is that for certain semisimple Lie groups $G$ of higher real rank, there also exist connected components of $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations, and which are nontrivial in the sense that they are not reduced to a single representation and its conjugates by $G$. The images in the $G$-character variety of these components are now called higher(-rank) Teichmüller spaces. We saw two examples in Section 3: